Please help me with those problems.
problem 1) If log 36 = a and log 125 = b
find log (1/12) in terms of a and b

problem 2) If log (base 4n) 40 square root of 3 = log (base 3n) 45
find n^3

Answer by stanbon(48535)

Problem 1: log 36 = log 6^2 = 2 log 6 = a
Therefore log 6 = a/2
log 125 = log 5^3 = 3 log5=b
Therefore log 5 = b/3
Now, log(1/12)= log(5/60) = log5-log60= log5-(log6+log10)
=log5-log6-log10
=b/3 = a/2 -1
Cheers,
Stan H.